OpenStudy (anonymous):

Consider the following cross: Parent 1: YyTt Parent 2: YyTt Using the rules of probability, determine the probability that the offspring will show YyTT genotype?

OpenStudy (anonymous):

Yy from Yy x Yy : the possibilities are YY, Yy, yY, yy : 2/4 give the right combination TT from Yy x Yy : the possibilities are YY, Yy, yY, yy : 1/4 give the right combination 2/4 times 1/4 = 2/16 = 1/8